If sin(a+2b)=root 3 /2 and cos(a-b)=1/2,find a and b.
Answer by lwsshak3(11628)

If sin(a+2b)=root 3 /2 and cos(a-b)=1/2,find a and b.
**
sin (a+2b)=√3/2
a+2b=60º
..
cos(a-b)=1/2
a-b=60º
..
a+2b=60º
a-b=60º
subtract
3b=0
b=0
a=60-b=60º